# How quantum and evolutionary algorithms can help each other: two examples

### v1

@@ Abstract
We investigate the potential of bio-inspired evolutionary algorithms for designing quantum circuits with specific goals, focusing on two particular tasks. The first one is motivated by the ideas of Artificial Life that are used to reproduce stochastic cellular automata with given rules. We test the robustness of quantum implementations of the cellular automata for different numbers of quantum gates The second task deals with the sampling of quantum circuits that generate highly entangled quantum states, which constitute an important resource for quantum computing. In particular, an evolutionary algorithm is employed to optimize circuits with respect to a fitness function defined with the Mayer-Wallach entanglement measure. We demonstrate that, by balancing the mutation rate between exploration and exploitation, we can find entangling quantum circuits for up to five qubits. We also discuss the trade-off between the number of gates in quantum circuits and the computational costs of finding the gate arrangements leading to a strongly entangled state. Our findings provide additional insight into the trade-off between the complexity of a circuit and its performance, which is an important factor in the design of quantum circuits.
@@
Quantum Computing, Quantum Circuits, Evolutionary Algorithms
@@ 1 Introduction
Quantum computing utilizes principles of quantum mechanics to perform certain computations faster and on/and larger scales than classical computers are capable of [1]. Quantum bits, or qubits, are key building blocks of quantum computers that enable the latter to be so advantageous in solving complex problems  [1]. The intersection of quantum computing and artificial intelligence (AI) is currently an area of active research that has the potential to revolutionize various fields and industry sectors. Quantum Machine Learning (ML) algorithms and, in particular, evolutionary algorithms, belong to one of the research directions that are actively explored now [2, 3, 4]. It is expected that the new results will help to speed up the optimization process further and improve the overall performance of computational algorithms. In particular, Evolutionary Quantum Computation (EQC) is an emerging approach that blends evolutionary algorithms and quantum computing to solve complex optimization problems [5, 6, 7].
@@
The aim of this work is twofold. First, we would like to show how quantum algorithms can be used to solve specific problems related to the engineering of evolutionary algorithms. Next, in a complimentary manner, we want to illustrate how evolutionary algorithms can be used to design quantum circuits fulfilling specific requirements. In particular, we will address the problem of designing quantum circuits that process initial product states into highly entangled states.
@@
Entanglement, one of the key concepts of quantum physics, also plays a crucial role in quantum information processing [8]. The term ’entangled state’ refers to a state of several quantum objects that are intrinsically correlated with each other, in a way different from all possible classical correlations [9]. F.e., when two qubits are entangled, changes in the state of one qubit affect the state of the other one. There are several measures proposed to quantify the strength of these essentially quantum correlations, and so a quantitative definition of the word ’highly,’ strictly speaking, depends on the chosen measure. Highly entangled quantum systems are usually difficult to simulate on classical computers [10, 11], and this fact has significant implications for quantum information processing technologies. Metaphorically speaking, highly entangled qubit states serve as fuel to quantum information processing and without them, the latter loses its superiority with respect to the classical processing [12]. To create a highly entangled state out of a non-entangled one is not an easy task. And even if we have managed to prepare a maximally entangled state, it can become less entangled just because the quantum system prepared in this state is interacting with the environment [13, 14, 15]. All real quantum systems suffer from this interaction whose action is conventionally described with the term "decoherence". Therefore, it is important to keep the time during which decoherence strikes the system as short as possible. Practically, this means to minimize the computation time. In addition, every time a quantum gate (that is a transformation of the state of a system of qubits) is implemented, an additional error occurs due to the non-ideal character of the implementation. After passing through many gates, the state is no longer reliable in the sense that it is very different from what it has to be according to the corresponding quantum program. Thus, the more complicated the quantum circuit is (in terms of the number of gates it consists of), the stronger the joint action of decoherence and accumulated implementation errors [16]. As has been mentioned above, not all entangled states are equal in their ’strength’. This means that we need to quantify and measure the amount of entanglement between different qubits. This brings another challenge, because, in a generic quantum system, the quantification of entanglement is a non-trivial task  [17].
Our study addresses two specific problems, which are discussed in Section 2. The first problem, addressed in  Subsection 2.1, involves realizing classical stochastic cellular automata using quantum circuits. This approach is innovative as it leverages quantum mechanics to enhance the modeling of complex systems, a topic of extensive interest in prior research [18]. The second problem, discussed in Subsection 2.2, focuses on designing quantum circuits that generate highly entangled multi-qubit states. This is, up to our knowledge, a new application of evolutionary algorithms: to optimize entanglement in quantum circuits. To this end we use the Meyer-Wallach entanglement measure as the fitness function. In Section 3 we summarize the results and discuss further perspective. Finally, in Section 4, we describe the methods used and highlight the relevant concepts.
@@ 2.1 Realizing stochastic cellular automata with quantum circuits
Initially, a predetermined number of chromosomes is generated, corresponding to the number of quantum gates per circuit. Through iterative refinement, the genetic algorithm evolves the existing chromosomes into new variations, singling out the fittest chromosome within each generation to serve as the parent for the subsequent generation. The assessment of the performance of the derived quantum circuits revolves around the measurement of the first qubit, $q_{0}$. This assessment entails comparing the measured probability of an initial state, computed from the population, with the corresponding target probability. To quantify the difference between two sets of probability distributions, we use the Kullback-Leibler divergence (KL) as a fitness function. The KL-fitness score measures the discrepancy between the distributions of the output states and the target states. In our case, these distributions are discrete, each corresponding to the eight different initial states of the three qubits. If the distributions completely match, the fitness function value is zero. A decreasing fitness value close to zero indicates an increase in the efficiency of the circuits.
The resulting quantum circuits are evaluated through measurements, thereby demonstrating the algorithm’s efficacy. For more details, see Supplementary Note I.
Neighbors Sto. CA Prob. Rule90 Rule110 Rand. Prob. 1 Rand. Prob. 2 Rand. Prob. 3
@@
Table 1: The deterministic and stochastic CAs considered in this paper. For stochastic cases, the values indicate the probability of an update of value 1 for the middle cells in the neighborhood of the triad. For the deterministic cases, the values indicate the exact update imposed.
We consider three different types of quantum CAs, starting with a stochastic critical cellular automaton, followed by more general stochastic CA (random updates) and a few deterministic rules, namely, rule $90$ and $110$  [19]. The updates for each type of CA are shown in Table 1. To enhance the robustness of our method, we limit the number of chromosomes by $N_{c}=20$ chromosomes and the number of generations by $N_{g}=500$. Simulations are then performed for $N_{ic}=50$ randomly chosen initial conditions. This approach builds upon our previous experiment, where we set $N_{c}=20$ chromosomes, $N_{g}=150$ generations, and $N_{ic}=20$ initial conditions were used [18].
Figure 1: (Left) The fitness scores as a function of the number of generations, for different number of gates. (Right) The number of gates vs. the best fitness scores. The fitness scores of each gate for the box plots are the best fitness scores per run and the fitness scores for the lower two plots are the average fitness scores of 50 runs.
Figure 2: (Left) The fitness scores as a function of the number of generations, for different mutation rates. (Right) Mutation rates vs. the best fitness scores. The fitness scores of each gate are the average fitness scores of 50 runs.
Figure 3: (Left) Fitness scores for different sets of probabilities against the number of generations for the KL-fitness function. (Right) Best fitness scores per run for KL-fitness function for different sets of probabilities. The fitness scores of each gate are the average fitness scores of 50 runs.
Figure 4: (Top) Comparison of the best fitness scores across all 50 runs over 500 generations for different CA rules. (Bottom) Visualization of the optimal quantum circuit: Achieving a fitness score of 0.294 using KL fitness function for Critical Stochastic Cellular Automaton. The circuit, composed of 15 gates with a 10% mutation probability, represents the most exceptional outcome within the study, underscoring the success of our approach in circuit optimization.
@@
Figure 2 illustrates the impact of different mutation rates on the KL-fitness score. The left panel of figure 2 shows that all mutation rates result in a decrease in fitness scores over generations, indicating improvement. However, mutation rates of 5%, 10%, and 20% stabilize more consistently compared to higher rates. The right panel box plots show that while higher mutation rates (30%, 50%) lead to decreased fitness scores, they introduce greater variability and less stability. These results highlight that while higher mutation rates can still improve circuit efficiency, the optimal balance between exploration and stability is achieved at a 20% mutation rate.
@@
We now test the above approach with generic deterministic and stochastic cellular CAs. We fixed parameters, by utilizing 15 gates and a 10% mutation probability. Specifically, we examine the application of our methodology to Rule 90 and Rule 110 (deterministic rules with fixed probabilities of 0 or 1 for the eight neighborhoods), as well as to eight randomly generated probabilities, each repeated three times, serving as target probabilities. The corresponding target probabilities for each experimental condition are detailed in Table 1, while the ensuing outcomes are visually represented in Figure 3. This comprehensive analysis showcases the versatility and effectiveness of our approach across a spectrum of rule types and probabilistic settings.
@@
The fitness scores obtained for the deterministic CAs, Rule 90 and Rule 110, are remarkably good, as shown in Figure 3 (left). This highlights the efficiency of our approach to the synthesis of quantum circuits for deterministic CAs. Notably, the outcomes stemming from the stochastic CA approach, characterized by the incorporation of randomly generated probabilities over three separate iterations, exhibit some promise. The optimal fitness scores attained for Random1, Random2, and Random3 are 0.3364, 0.1681, and 0.0014, respectively. It is worth mentioning that while Random1 achieves a commendable fitness score of 0.3364, indicating its promising performance, there is potential for further refinement of the evolutionary process to reduce any existing disparities. A comprehensive evaluation of the fitness scores across different probability sets, as presented in Figure 3 (right), to the number of generations, provides valuable insights into the performance trends for the deterministic CA models (Rule 90 and Rule 110), as well as the stochastic CA model utilizing randomly generated probabilities over three iterations (Random1, Random2, and Random3), all conducted with an initial set of 50 run conditions.
@@
Figure 4 (Left) illustrates the fitness scores corresponding to the diverse CA rules. The best fitness score of 0.294 was achieved by the critical CA. Notably, for two deterministic CA rules, namely Rule 90 and Rule 110, the fitness scores approximate zero, reflecting a deliberate pursuit of minimizing the absolute disparity between initial and final states. This trend aligns logically with the deterministic nature of these CA rules, where probability values exclusively assume binary states of 0 or 1. In contrast, the scenario changes when dealing with randomly generated stochastic CA rules. In this context, the fitness scores demonstrate variability across three distinct randomly assigned probabilities. Intriguingly, among these variations, Random 3 emerges as the front-runner, boasting the most favorable fitness scores. Remarkably, all three sets of randomly generated probabilities exhibit fitness values superior to that of the critical stochastic CA value. Nevertheless, to unravel the underlying mechanisms driving these observed phenomena, a comprehensive exploration beckons. This entails conducting a more extensive array of experiments, encompassing increased generations and a higher replication count. The intricate interplay of factors contributing to fitness score disparities between distinct randomly generated probabilities necessitates a more exhaustive investigation.
@@
In Figure 4 (Right), we show the circuit that significantly outperforms the average outcomes of runs associated with the critical stochastic cellular automaton. Impressively, this particular realization achieved a fitness score of 0.294, as evaluated by the KL fitness function outlined in Equation (1).

### v2

# Optimizing quantum circuits with evolutionary algorithms for stable Boolean gates, elementary cellular automata, and highly entangled quantum states

@@ Abstract
We investigate the potential of bio-inspired evolutionary algorithms (EAs) for designing quantum circuits with specific goals, focusing on two particular tasks. The first task involves using these algorithms to reproduce stochastic cellular automata with given rules. We test the robustness of quantum implementations of the cellular automata for different numbers of quantum gates. The second task deals with sampling quantum circuits that generate highly entangled quantum states, which constitute an important resource for quantum computing. In particular, an evolutionary algorithm is employed to optimize circuits with respect to a fitness function defined with the Meyer-Wallach (MW) entanglement measure. We demonstrate that, by balancing the mutation rate between exploration and exploitation, we can find entangling quantum circuits for up to five qubits. We also discuss the trade-off between the number of gates in quantum circuits and the computational costs of finding the gate arrangements leading to a strongly entangled state. Our findings provide additional insight into the trade-off between the complexity of a circuit and its performance, which is an important factor in the design of quantum circuits.
@@
Quantum Computing, Quantum circuits, Quantum Gates, Entanglement, Evolutionary Algorithms
@@ 1 Introduction
Quantum computing utilizes principles of quantum mechanics to perform certain computations faster and on larger scales than classical computers are capable of [41]. Quantum bits, or qubits, are key building blocks of quantum computers that enable the latter to be advantageous in solving complex problems [41]. The intersection of quantum computing and artificial intelligence (AI) is currently an area of active research that has the potential to revolutionize various fields and industry sectors. Quantum Machine Learning (QML) algorithms, particularly EAs, belong to one of the research directions actively explored now [14, 45, 54]. It is expected that the new results will help to speed up the optimization process further and improve the overall performance of computational algorithms. In particular, Evolutionary Quantum Computation (EQC) is an emerging approach that blends EAs and quantum computing to solve complex optimization problems [51, 16, 7], with recent applications including quantum circuit optimization [52], quantum state engineering [16], and hybrid quantum genetic algorithms for optimization tasks [8]. It is expected that these advancements will help to speed up the optimization process and improve the overall performance of computational algorithms.
@@
The aim of this work is twofold. First, we would like to show how quantum algorithms can be used to solve specific problems related to the engineering of EAs. Next, in a complementary manner, we want to illustrate how EAs can be used to design quantum circuits fulfilling specific requirements. In particular, we will address the problem of designing quantum circuits that process the initial product states into highly entangled states.
@@
Entanglement, one of the key concepts of quantum physics, also plays a crucial role in quantum information processing [60]. The term ‘entangled state’ refers to a state of several quantum objects that are intrinsically correlated with each other, in a way different from all possible classical correlations [28]. For example, when two qubits are entangled, changes in the state of one qubit affect the state of the other one. There are several measures proposed to quantify the strength of these essentially quantum correlations, and so a quantitative definition of the word ‘highly,’ strictly speaking, depends on the chosen measure. Highly entangled quantum systems are often difficult to simulate on classical computers[57, 56], though certain exceptions, such as states generated by Clifford circuits, can be efficiently simulated[3]. This complexity has significant implications for quantum information processing technologies. Metaphorically speaking, highly entangled qubit states serve as fuel to quantum information processing, and without them, the latter loses its superiority with respect to classical processing [17]. To create a highly entangled state out of a non-entangled one is not an easy task. And even if we have managed to prepare a maximally entangled state, it can become less entangled just because the quantum system prepared in this state is interacting with the environment [47, 2, 4]. All real quantum systems suffer from this interaction, whose action is conventionally described with the term “decoherence”. Therefore, it is important to keep the time during which decoherence strikes the system as short as possible. Practically, this means minimizing the computation time. In addition, every time a quantum gate (that is, a transformation of the state of a system of qubits) is implemented, an additional error occurs due to the non-ideal character of the implementation. After passing through many gates, the state is no longer reliable in the sense that it is very different from what it has to be according to the corresponding quantum program. Thus, the more complicated the quantum circuit is (in terms of the number of gates it consists of), the stronger the joint action of decoherence and accumulated implementation errors [50]. As has been mentioned above, not all entangled states are equal in their ‘strength’. This means that we need to quantify and measure the amount of entanglement between different qubits. This brings another challenge, because, in a generic quantum system, the quantification of entanglement is a non-trivial task [23].
In Section 2 we outline the methodology and introduce the key concepts. Section 3 then presents two main results. Subsection 3.1 shows how classical stochastic cellular automata can be implemented with quantum circuits, extending earlier work on complex-system modeling [13]. Subsection 3.2 introduces a novel use of evolutionary algorithms to design circuits that maximize multi-qubit entanglement, evaluating each candidate with the Meyer–Wallach measure. Finally, Section 4 summarizes the findings and outlines future directions.
2 Methods
Our study implements two distinct methods for designing quantum circuits. The first method utilizes cellular automata rules combined with the KL fitness function. Here we advanced the original framework from Ref. [13] for generating quantum circuits that replicate quantum circuits with EAs. Upon this groundwork, the second method focuses on designing quantum circuits for maximum entanglement, using the MW measure and von Neumann entropy for entanglement evaluation. The comprehensive details regarding the implementation of evolutionary quantum algorithms for both methodologies are thoroughly delineated in Supplementary Note I.
2.1 Evolutionary algorithms: Mutations and fitness function
The evolutionary algorithm initiates with a specific assigned number of chromosomes, each populated with a set number of quantum gates within the circuit. This initial population is then subject to evolution, where the fittest chromosomes are identified, iteratively refining the existing ones, and introducing the best-fit chromosome at the forefront of each new generation. This evolutionary cycle continued for a predetermined number of generations, with each generation’s best chromosome becoming the parent for the next. Throughout this process, the chromosomes are subjected to mutations to enhance optimization. The mutation process is introduced in two different forms: either by swapping existing gates in the chromosome with randomly selected ones from the gate pool, or by entirely replacing the chromosome to identify the best four parental candidates suitable for the next generation. The mutation process is controlled by a set of probabilities. The four best chromosomes are preserved as "elites" without changes, whereas the remaining chromosomes are selected for evolution based on their fitness, with selection changes directly linked to their performance.
To evaluate the performance of the quantum circuits, we measure only the first qubit (q0) and consider the fitness function as the $D_{KL}$, which measures the difference between two probability distributions and has been used in other works [34, 33] as a fitness function:
2.2 Assessing entanglement in basic quantum states of qubits
[…]
In this state, when measuring one qubit, there is no certainty about the state of the other qubit. Namely, there is a probability of $1/2$ that the other qubit is in state 0 and a probability of $1/2$ that it is in state $1$.
In general, suppose we have two quantum systems $Q_{1}$ and $Q_{2}$. These systems are said to be entangled if their joint quantum state cannot be factored into a product of individual states. More formally, a two-qubit state $|\psi\rangle$ is entangled if it cannot be written as the tensor product of two pure states, i.e., $|\psi\rangle\neq|\psi_{1}\rangle\otimes|\psi_{2}\rangle$ , where $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$ are states belonging to individual systems $Q_{1}$ and $Q_{2}$, respectively. This condition holds for pure states, and for mixed states, the density matrix of the system cannot be written as a convex combination of product states. The simplest examples of entangled states are the Bell states, which represent maximally entangled two-qubit states [58], and they enable tasks such as quantum cryptography [24], superdense coding [29], teleportation [11], and entanglement swapping [38]. These Bell states are:
$|\Phi^{+}\rangle$ $=\frac{1}{\sqrt{2}}\left(|0\rangle\otimes|0\rangle+|1\rangle\otimes|1\rangle\right),\quad|\Phi^{-}\rangle=\frac{1}{\sqrt{2}}\left(|0\rangle\otimes|0\rangle-|1\rangle\otimes|1\rangle\right),$
$|\Psi^{+}\rangle$ $=\frac{1}{\sqrt{2}}\left(|0\rangle\otimes|1\rangle+|1\rangle\otimes|0\rangle\right),\quad|\Psi^{-}\rangle=\frac{1}{\sqrt{2}}\left(|0\rangle\otimes|1\rangle-|1\rangle\otimes|0\rangle\right)$
where $|00\rangle,|01\rangle,|10\rangle,|11\rangle$ are the computational basis states.
For example, the measure of expressibility can be calculated by taking the $D_{KL}$ between the estimated fidelity distribution and that of the Haar-distributed ensemble [30]. In this paper, we considered two measures, the MW [36] and the entanglement entropy (EE)[19, 40, 32]. The MW entanglement measure is popular because of its scalability and ease of computation. It considers a pure state of $n$ qubits of the form [15]:
[…]
Since the purity of the state of qubit $k$ is given by $Tr[\rho_{k}^{2}]=\braket{\hat{x}^{k}|\hat{x}^{k}}^{2}+\braket{\hat{y}^{k}|\hat{y}^{k}}^{2}$, we obtain [15]
[…]
which is the usual definition of the MW measure. A detailed derivation and discussion of this measure, especially its implications for both pure and mixed states, are provided in Supplementary Note II.
$S(\rho_{A})=-\mathrm{Tr}(\rho_{A}\log_{2}\rho_{A})\,$, (10)
In what concerns implementation, we have used the mathematical formulations presented in Equations (9) and (10) in a framework previously developed, aligned with the methodologies outlined in Ref. [13]. The EAs have emerged as a promising tool for quantum circuit design [53, 16, 7]. So our implementation of EAs for quantum circuit design aims to maximize entanglement using two specified methods. The framework employs an evolutionary algorithm to automatically generate quantum circuits that satisfy defined properties specified through a fitness function. The evolution properties can be controlled by tuning parameters such as initial population, number of generations, probability of mutation operator, number of gates in the circuit, and the chosen fitness function. Details of the implementation process are available in Supplementary Note III.
@@ 3.1 Realizing stochastic cellular automata with quantum circuits
Initially, a predetermined number of chromosomes is generated, corresponding to the number of quantum gates per circuit. Through iterative refinement, the genetic algorithm evolves the existing chromosomes into new variations, selecting the fittest chromosome within each generation to serve as the parent for the subsequent generation. The assessment of the performance of the derived quantum circuits revolves around the measurement of the first qubit, $q_{0}$. In our quantum circuits, we denote qubits as $q_{i}$, where $i$ indexes the qubit position (e.g., $q_{0}$ is the first qubit in a three-qubit system). This assessment entails comparing the measured probability of an initial state, computed from the population, with the corresponding target probability. To quantify the difference between two sets of probability distributions, we use the Kullback-Leibler divergence ($D_{KL}$) [9] as a fitness function. The $D_{KL}$ fitness score measures the difference between two probability distributions, with a detailed formulation provided in Section 2.1, (Equation 1). In our case, these distributions are discrete, each corresponding to the eight different initial states of the three qubits. If the distributions completely match, the fitness function value is zero. A decreasing fitness value close to zero indicates an increase in the efficiency of the circuits. The resulting quantum circuits are evaluated through measurements, thereby demonstrating the algorithm’s efficacy. For more details, see Supplementary Note I.
Neighbors Sto. CAProb. Rule90 Rule110 Prob. 1 Prob. 2 Prob. 3
@@
Table 1: The deterministic and stochastic CAs are considered in this paper. For stochastic cases, the values indicate the probability of an update of value 1 for the middle cells in the neighborhood of the triad. For the deterministic cases, the values indicate the exact update imposed.
We consider three different types of CAs, starting with a stochastic critical cellular automaton, followed by more general stochastic CA (random updates) and a few deterministic rules, namely, rule $90$ and $110$ [59]. The updates for each type of CA are shown in Table 1. Our evolutionary algorithm uses a population of $N_{c}$ chromosomes (where $N_{c}$ is the number of candidate circuits), evolves over $N_{g}$ generations (where $N_{g}$ is the number of iterations), and tests $N_{ic}$ initial conditions (where $N_{ic}$ is the number of random starting states). In this study, we set $N_{c}=20$, $N_{g}=500$, and $N_{ic}=50$ to enhance robustness, building on prior settings of $N_{c}=20$, $N_{g}=150$, and $N_{ic}=20$ [13]. Recent advancements in quantum cellular automata (QCA) models further contextualize our work. For instance, Arrighi et al. (2022) [6] demonstrated that QCA can simulate complex lattice gauge theories, suggesting a quantum analog to our classical CA replication, while Jones et al. (2022) [31] experimentally realized a QCA on a 23-qubit digital superconducting processor, achieving robust rule replication under noise, paralleling our focus on stability across gate counts (Figure 1).
a)Refer to captionb)
Figure 1: a) The fitness scores as a function of the number of generations, for different number of gates. b) The number of gates vs. the best fitness scores. The fitness scores of each gate for the box plots are the best fitness scores per run, and the fitness scores for the lower two plots are the average fitness scores of 50 runs.
a)Refer to captionb)
Figure 2: a) The fitness scores as a function of the number of generations, for different mutation rates. b) Mutation rates vs. the best fitness scores. The fitness scores of each gate are the average fitness scores of 50 runs.
a)Refer to captionb)
Figure 3: a) Fitness scores for different sets of probabilities against the number of generations for the KL-fitness function. b) Best fitness scores per run for $D_{KL}$ fitness function for different sets of probabilities. The fitness scores of each gate are the average fitness scores of 50 runs.
a)
Refer to captionb)
Figure 4: a) Comparison of the best fitness scores across all 50 runs over 500 generations for different CA rules. b) Visualization of the optimal quantum circuit: Achieving a fitness score of 0.294 using the $D_{KL}$ fitness function for Critical Stochastic Cellular Automaton. The circuit, composed of 15 gates with a 10% mutation probability, represents the most exceptional outcome within the study, underscoring the success of our approach in circuit optimization.
@@
Figure 2 illustrates the impact of different mutation rates on the $D_{KL}$ fitness score. Figure 2 a) shows that all mutation rates result in a decrease in fitness scores over generations, indicating improvement. However, mutation rates of 5%, 10%, and 20% stabilize more consistently compared to higher rates. Figure 2 b) shows that while higher mutation rates (30%, 50%) lead to decreased fitness scores, they introduce greater variability and less stability. These results highlight that while higher mutation rates can still improve circuit efficiency, the optimal balance between exploration and stability is achieved at a 20% mutation rate.
@@
We now test the above approach with generic deterministic and stochastic cellular CAs. We fixed parameters by utilizing 15 gates and a 10% mutation probability. Specifically, we examine the application of our methodology to Rule 90 and Rule 110 (deterministic rules with fixed probabilities of 0 or 1 for the eight neighborhoods), as well as to eight randomly generated probabilities, each repeated three times, serving as target probabilities. The corresponding target probabilities for each experimental condition are detailed in Table 1, while the ensuing outcomes are visually represented in Figure 3. This comprehensive analysis showcases the versatility and effectiveness of our approach across a spectrum of rule types and probabilistic settings.
@@
The fitness scores obtained for the deterministic CAs, Rule 90 and Rule 110, are remarkably good, as shown in Figure 3 a). This highlights the efficiency of our approach to the synthesis of quantum circuits for deterministic CAs. Notably, the outcomes stemming from the stochastic CA approach, characterized by the incorporation of randomly generated probabilities over three separate iterations, exhibit some promise. The optimal fitness scores attained for Random1, Random2, and Random3 are 0.3364, 0.1681, and 0.0014, respectively. It is worth mentioning that while Random1 achieves a commendable fitness score of 0.3364, indicating its promising performance, there is potential for further refinement of the evolutionary process to reduce any existing disparities. A comprehensive evaluation of the fitness scores across different probability sets, as presented in Figure 3 b), to the number of generations, provides valuable insights into the performance trends for the deterministic CA models (Rule 90 and Rule 110), as well as the stochastic CA model utilizing randomly generated probabilities over three iterations (Random1, Random2, and Random3), all conducted with an initial set of 50 run conditions.
@@
Figure 4 a) illustrates the fitness scores corresponding to the diverse CA rules. The best fitness score of 0.294 was achieved by the critical CA. Notably, for two deterministic CA rules, namely Rule 90 and Rule 110, the fitness scores approximate zero, reflecting a deliberate pursuit of minimizing the absolute disparity between initial and final states. This trend aligns logically with the deterministic nature of these CA rules, where probability values exclusively assume binary states of 0 or 1. In contrast, the scenario changes when dealing with randomly generated stochastic CA rules. In this context, the fitness scores demonstrate variability across three distinct randomly assigned probabilities. Intriguingly, among these variations, Random 3 emerges as the front-runner, boasting the most favorable fitness scores. Remarkably, all three sets of randomly generated probabilities exhibit fitness values superior to that of the critical stochastic CA value. Nevertheless, to unravel the underlying mechanisms driving these observed phenomena, a comprehensive exploration beckons. This entails conducting a more extensive array of experiments, encompassing increased generations and a higher replication count. The intricate interplay of factors contributing to fitness score disparities between distinct randomly generated probabilities necessitates a more exhaustive investigation.
@@
In Figure 4 b), we show the circuit that significantly outperforms the average outcomes of runs associated with the critical stochastic cellular automaton. Impressively, this particular realization achieved a fitness score of 0.294, as evaluated by the $D_{KL}$ fitness function outlined in Equation (1).